OpenStudy (anonymous):

Simplify the radical expression. the square root of 72x^2

OpenStudy (tkhunny):

What do you think?

OpenStudy (anonymous):

\[36\sqrt{x ^{2}}\]?

OpenStudy (tkhunny):

You divided by 2. Why would that be right? You need to find perfect squares and remove them. 72 = 2*36 = \(2^{2}\cdot 18 = 2^{2}\cdot 3^{2}\cdot 2\) So, 72 has perfect squares of both 2 and 3.

OpenStudy (anonymous):

Okay, I think I understand now! thanks!

OpenStudy (tkhunny):

The \(x^{2}\)-part is a little tricky.

OpenStudy (whpalmer4):

Yes, indeed — you can't blindly replace \(\sqrt{x^2}\) with \(x\) unless you somehow know that \(x>0\)